From GlueXWiki
Jump to: navigation, search

November 15, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • First package testing (Vlad)
    • Plans for finishing/testing packages #1 and #4
  2. Corrosion tests (Vlad)
  3. Engineering (Bill)
  4. Electronics (Chris, Nick)
  5. Third package tests: wire resolution FDC E-log (Lubomir)
    • Moving to Blue Crab?
  6. Other


Participants: Eugene, Bill, Dave, Chris, Nick, Vlad, Simon, Beni, and Lubomir.


- Tina finished wire soldering of wire plane #30. Tension measurements will be done today. Tomorrow Chris will solder the components. Dave will deaden the wires either on Friday or Monday. So, the plane will be ready for installation in package #4 on Tuesday.

- Vlad finished testing package #1. All channels are working fine. Three pre-amps are oscillating, but the channels are working if using other cards at the same places. Two of the cards have been replaced already with new ones, still the same problem. After ground installation we hope the problem will disappear, if not we can replace the cards later.

- Package #1 was disassembled from the mounting table, we still have three metal pins in the package. It was moved to granite table and Vlad just finished the installation of the grounding clips. Next step (probably tomorrow) will be the installation of the cooling tubes. Bill made a tool to place the tube ends in the proper positions and cut them, then install the plastic manifolds. He explained how to fix the tubes during this procedure so that they don't apply pressure on the cards/cathodes.

- We plan to move the third package back to Blue Crab on Monday. We will take from it the strong-back, needed for testing the package #1 at 126, and the hubs needed to finish the installation of package #4. On Tuesday we plan to install the last two cells of package #4. It will be flushed with gas over Thanksgiving and Vlad will test it the week after Thanksgiving. Package #1 will be moved to 126 for testing the first week of December.

- 6 sets of cathode foils have been shipped on Monday. Probably they will not be able to make 18 sets with the material they have. Bill found that the 2micron-copper/Kapton foil is still available. Eugene suggested to buy material for 14 sets. Bill will ask for a quotation for 14 and 24 sets.

Cathode corrosion test

- Test #3 is going on: Ar/CO2 gas, at ~70degC. This is to study the activation energy, hopefully will have some results by Christmas. Vlad will take cathode and O-ring samples and will continue the tests at Moscow. He's leaving November 30.


- Bill: the new SMC chiller and the fluorinert will come middle of December, finished the documentation for the pressure system, discussed with Ron Bartek the power for the chiller in 126. The chiller requires also colling with water, still not clear how to do this.

- AC at 126: As Eugene suggested it's not worth investing in Blue Crab, so we will do the tests at 126. Before it was proposed to install two ceiling fans, but then a cheaper solution was found: two split cooling units. Bill after the meeting: the reqs for the units are out, expect to install them before Christmas.

- FDC installation. The FDC installation cart is already at Blue Crab. First, it has to surveyed and the rail adjusted. All the packages will be installed on the cart at Blue Crab and surveyed there. We don't know when we will finish the complete tests of the packages; it depends on the fADC125 availability. Eugene: we can bring the FDC to the Hall middle of July 2013 and will have time for installation/testing till the end of October. What has to be done there: once more FDC survey (few days), cable installation (two weeks at least excluding testing of the connectivity), installation of/connection to the gas, cooling, HV, LV systems (two weeks), testing/troubleshooting/ground improvements (at least a month, maybe two). Will need to have all the electronics, DAQ system working by that time. Will have to install several scintillator pedals for cosmic triggering.


- Chris: no news about the daughter cards, it was three days delivery and we are waiting already a month. Will get the grounding clips right after Thanksgiving.

Third package tests

- As explained in FDC-Elog #96, the wire resolution was estimated. In addition to the wire TDC information, six cathode cards were used on cells #1,3, and 4. Using only cathode information from cells #1 and #4 that have parallel wires and strips, one finds a line on cell #3. Using the cathode information from cell #3 defines another line and their crossing gives you the hit point (even the angle is not completely defined) and the distance to the nearest wire can be calculated. This distance is compared with the distance obtained from the timing information for that wire using some initial time-to-distance conversion. The difference between to the distances has sigma of ~320microns for drift times between 30 and 160 ns. After correcting for the cathode resolution one obtains ~190 microns wire resolution, finally! This method can be used to extract the time-to-distance function experimentally.

- There's still some noise on cells #1 and #2. Beni thinks it requires better grounding. Chris will make additional ground that will connect to the fast-ons on the wire frames. After the meeting Beni and Lubomir looked at the noise. Swapping the translation boards didn't affect the noise position, however swapping the cables at the chamber also didn't change the noise. Therefore, the main suspects are the TDCs or a combination of the TDCs and the other elements. Beni will look again tomorrow.


- No FDC meetings next two weeks: on Thanksgiving and the week after that. Will meet again on December 6.